1)On the first day of school, the percentage of boys in a particular class is 60%. During the school year, six girls move away, and are replaced in the class by six boys; this makes the class roster 75% boys. Find the number of boys and girls in the class on the first day of school.


2)An express and local train leave GraysLake at 3 P.M. and head for Chicago 50 miles away. The express travels twice as fast as the local and arrives 1 hour ahead of the local. Find the speed of each train.

Answer by stanbon(75887)

On the first day of school, the percentage of boys in a particular class is 60%.
During the school year, six girls move away, and are replaced in the class by six boys; this makes the class roster 75% boys.
Find the number of boys and girls in the class on the first day of school.
----------
Original class numbers:
Total = "x"
Boys = 0.6x
Girls= 0.4x
------------------
After 6-girls move and are replace by 6 boys:
Total = "x"
Boys:= 0.6x+6 = 0.75x
--------------------------
Boys equation shows:
0.6x + 6 = 0.75x
6 = 0.15x
x = 40 (number in the class on the 1st day)
-----------------
2)An express and local train leave GraysLake at 3 P.M. and head for Chicago 50 miles away.
The express travels twice as fast as the local and arrives 1 hour ahead of the local. Find the speed of each train.
-----------
Express DATA:
Distance = 50 miles : Rate = 2x mph ; time = 50/2x
---------
Local train DATA:
Distance = 50 miles ; Rate = x mph ; time = 50/x
-----------------
EQUATION:
local train time - express train time = 1 hr
50/x - 50/2x = 1
Multiply thru by 2x to get:
100 - 50 = 2x
50 = 2x
x = 25 mph ( speed of the local train)
2x = 50 mph ( speed of the express)
=================
Cheers,
Stan H.